OpenStudy (anonymous):

Which is a true statement about the hypotenuse of a right triangle? It is the shortest side. It meets a leg to form a right angle. Its length is equal to the sum of the length of the two legs. It lies across from the right angle.

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (anonymous):

its 3 or 4....

OpenStudy (anonymous):

or its C or D

OpenStudy (phi):

do you know the pythagorean theorem? a^2 + b^2 = c^2 where a and b are the legs and c is the hypotenuse

OpenStudy (phi):

you are suppose to think, if a^2 + b^2 = c^2 is true, then a+b= c is not true

OpenStudy (anonymous):

wait it is C or D correct?

OpenStudy (anonymous):

Its D